how do i find the length of the radius if the circumference 10 feet?

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

how do i find the length of the radius if the circumference 10 feet?
---
circumference = (pi)d
---
10 = (pi)d
----
d = 10/(pi) feet
======================
radius = d/2 = 5/pi feet
========================
Cheers,
Stan H.
===============

Answer by solver91311(24713)





so



John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism